OpenStudy (anonymous):

If 2x+2x+2x=12, what is the value of 2x-1?

OpenStudy (anonymous):

hello just solve for X then replace in 2x - 1

OpenStudy (anonymous):

wait i still dont get it

OpenStudy (anonymous):

example 3x + 2x = 10 5x = 10 x = 10/5 x = 2

OpenStudy (anonymous):

Oh I see you just combined 3x and 2x to make 5 and then divided 10 by 5 which gives you 2. But im just confused by the part where you have to replace the answer in 2x-1

OpenStudy (anonymous):

Because apparently the answer is 3 but i dont get how its is since i keep getting 6

OpenStudy (anonymous):

for eg you need 5x - 3 and you got x = 5 then you do 5 ( 5 ) - 3 25 - 3 = 22 finally you say 5x - 3 = 22

OpenStudy (anonymous):

come try it

OpenStudy (anonymous):

then 5x - 3=22 +3 +3 5x= 26 5 = 5 x=5.2 am i right??